# Optimal Estimation of Temperature

### v4

@@ Abstract
I Introduction
Temperature fluctuations become increasingly apparent as the system size decreases. The significant fluctuations of temperature in finite-sized systems challenge the zeroth law of thermodynamics and the notion of a well-defined temperature for the finite-sized system. Such challenges have sparked a wide range of subsequent studies, including the definition of entropy and temperature in isolated systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], the existence of negative absolute temperatures [11, 12, 13, 14, 15, 16, 17, 18, 19, 20], high resolution thermometer (HRT) [21, 22, 23, 24], quantum calorimeter [25] or quantum thermometry [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], the energy–temperature uncertainty relation (ETU) [38, 39], the thermodynamics of finite-sized systems [40, 41, 42, 43, 44, 45], and nonequilibrium processes and stochastic thermodynamics [46, 47, 48, 49]. Despite considerable progress, conclusions drawn from specific viewpoints or scenarios are sometimes inconsistent, and a unified understanding of thermal fluctuations in finite-sized systems remains incomplete. This situation reflects the absence of a general mathematical framework for studying such fluctuations, which motivates our work to develop a more systematic approach.
Historically, Einstein firstly investigated equilibrium fluctuations by relating the probability of macroscopic states to entropy through an inversion of the Boltzmann’s entropy formula [50, 51, 52, 53]. Bohr suggested that there should exist a form of complementarity between temperature and energy in thermodynamics similar to that of position and momentum in quantum theory [38, 39]. Then, Mandelbrot linked the ETU ($\beta$ denotes the inverse temperature of the system and $E$ denotes the energy of the system)
@@ I Introduction
to the Cramér–Rao bound in estimation theory through the polemic exchange with Feshbach and Kittel [54, 55, 56, 57], where $\Delta(\cdot)=\sqrt{\langle\cdot^{2}\rangle-\langle\cdot\rangle^{2}}$. This practice originated from 1956 [58] when Mandelbrot applied the estimation theory to the derivation of statistical thermodynamics from purely phenomenological principles (also see [59, 60, 61, 62]). Experimentally, Gaussian-type temperature fluctuations have been observed in paramagnetic salt HRT [21, 22, 23], ferromagnetic alloy HRT [24], and a nanoscale electron calorimeter [25]. They verified that the temperature of the system exhibits fluctuations, while the temperature of the reservoir remains fixed.
@@
In repeated sampling, the optimal estimation of temperature exhibits sample-size dependence, coinciding with their counterparts in nanothermodynamics [40, 41, 42, 43, 44, 45]. Actually, the “sampling ensemble” can be regarded as the statistical interpretation of the replica trick in Hill’s nanothermodynamics. And a Gaussian-type temperature fluctuation naturally emerges from a general temperature distribution due to the central limit theorem (CLT).
Compared with quantum thermometry, which mainly focuses on measuring temperature via quantum coherence in few-body systems [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], our work provides a framework for describing temperature fluctuations in finite-sized systems, offering an alternative approach to achieve more accurate temperature estimates in many-body systems.
II Estimation of temperature in single sampling
Developing Mandelbrot’s viewpoint about temperature fluctuations [56, 58, 59, 60], i.e., “temperature for systems-in-isolation should be viewed as a statistical estimate of the parameter of a conjectural canonical distribution”, we treat a finite-sized system in canonical ensemble as a thermometer of a reservoir. The $N$-particle system, with $i\in D$ denoting its microstate, contacts a real or hypothetical heat reservoir with temperature $T$, where $D$ denotes the set of energy level indices. At equilibrium, the probability of the system at state $i$ obeys the Boltzmann-Gibbs formula:
@@ II Estimation of temperature in single sampling
where $\beta=T^{-1}$ denotes the inverse temperature, $E_{i}$ $(d_{i})$ denotes the energy (degeneracy) of state $i$ and $Z=\sum_{i\in D}d_{i}e^{-\beta E_{i}}$ is the partition function.
@@
where $\sigma(E)\equiv\sum_{i\in D}\delta(E-E_{i})d_{i}$ denotes the density of states , $\text{d}_{E}^{0}\sigma(E)\equiv\sigma(E)$ and [65]
$\text{d}^{k}_{E}\sigma(E)\equiv\begin{cases}\frac{\text{d}^{k}\sigma(E)}{\text{d}E^{k}}&\text{for }k\textgreater 0\\ \int_{E_{\text{min}}}^{E}\text{d}E^{{}^{\prime}}\text{d}_{E^{\prime}}^{k+1}\sigma(E^{\prime})&\text{for }k\textless 0\end{cases}.$ (7)
In the derivation, we have assumed the density of states $\sigma(E)$ satisfies
@@
for integer $l\in[0,k)$, which holds in classical systems but not in quantum systems. This is relevant to the third law of thermodynamics [66].
When the above assumption (8) is invalid, the unbiased estimator of $\beta^{k}$ for $k<0$, as well as that of $\beta$ (except for an exponentially decaying bias with $N$), are still given in Eq. (6). In FIG.1, we illustrate this bias in an analogous low-temperature Fermi gas system using relative bias $\Phi_{\beta}=|(\langle\hat{\beta}^{(1)}\rangle-\beta)/\beta|$.
@@
It is worth noting that $k$ has an upper bound (in this paper we mainly consider the case $k=\pm 1$), depending on the degrees of freedom of the system. As a example, taking the density of states $\sigma(E)=E^{3N/2-1}$ of a classical ideal gas as an example [67], $\hat{\beta}^{(k)}$ in Eq. (6) is
@@
where $k\leq 3N/2-2$ and $\Gamma$ denotes the Gamma function [68].
@@
Secondly, when $E_{\text{max}}$ is finite, the unbiased estimator of inverse temperature is still given in Eq.(6), regardless of whether the inverse temperature is positive or negative. However, it exhibits an exponentially decaying bias with increasing $N$. In FIG.2, we illustrate this bias in an $N$ non-interacting two-level systems, showing that it leads to a relative bias $\Phi_{\beta}\sim\text{e}^{-\alpha_{1}N}$, where the decay rate $\alpha_{1}=\ln(1+\text{e}^{-|\beta|\varepsilon})$ (see Appendix C), and $\varepsilon$ denotes the spacing of energy levels.
For positive temperatures, the unbiased estimator of temperature remains given by Eq. (6), except for a similar exponentially decaying bias. In contrast, for negative temperatures, using Eq. (6) to estimate the system’s temperature results in a bias that grows exponentially, which demonstrates that Gibbs temperature cannot be used to estimate the temperature of a negative-temperature system. In FIG.3, we illustrate these biases in an $N$ non-interacting two-level systems, showing that they lead to a relative bias $\Phi_{T}=|(\langle\hat{\beta}^{(-1)}\rangle-T)/T|\sim\text{e}^{-\alpha_{2}N}$, where the decay (increase) rate $\alpha_{2}=\frac{\beta\varepsilon}{2}+\ln\cosh(\frac{\beta\varepsilon}{2})$ (see Appendix C). The above indicates that an optimal estimator for the temperature of a negative-temperature system likely does not exist. Moreover, we extend the unbiased estimator of $\beta^{\alpha}$ to $\alpha\in\mathbb{R}$ (see Appendix D).
Figure 3: Relative bias of $T$ varies as NN when $E_{\text{max}}$ is finite. Here, the density of states is the same as FIG.2 (see Appendix C). (a) $T>0$. (b) $T<0$.
@@
A theorem [63, 70, 71] following the RBLS theorem is
Theorem 2.
Let $T$ be a complete sufficient statistic for a parameter $\theta$, and let $\tilde{\theta}(T)$ be any estimator based only on $T$. Then $\tilde{\theta}(T)$ is the unique UMVUE of its expected value.
Theorem 2 tells us that the UMVUE of temperature in the canonical ensemble is a function of its complete sufficient statistic. Meanwhile the Boltzmann–Gibbs formula in Eq. (5) belongs to the exponential family in statistical inference, and its complete sufficient statistic is $E_{i}$ [63]. Because the estimators $\hat{\beta}^{(k)}$ in Eq. (6) are not only unbiased but also a function of $E_{i}$, the estimators of $\hat{\beta}^{(k)}$ are the UMVUEs of $\beta^{k}$ according to Theorem 2. Moreover, no simple functional relationship exists among the UMVUEs of different moments of $\beta$, indicating that a unique definition of temperature does not exist in finite-sized systems.
@@
Hence, the Boltzmann entropy $S_{B}$ and the Gibbs entropy $S_{G}$ correspond to the UMVUE of inverse temperature and temperature respectively. It is worth mentioning that $\hat{\beta}_{B}$ and $\hat{T}_{G}$ have been discussed in previous studies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. The starting points of these studies are almost entirely based on the four laws of thermodynamics, which are established on empirical facts about systems in the thermodynamic limit. However, for finite-sized systems, whether these empirical facts and the four laws of thermodynamics are still valid is doubtful and unconfirmed. In contrast, we base our approach on the mathematical theorems in statistical inference, which is rigorous.
@@
    Determine the expressions for $\sigma$ and $\Omega$, which are either known before the measurement or can be obtained using the method in Section III. This method relates $\sigma$ to the partition function and its derivatives in the large-$N$ limit, allowing us to calculate $\sigma$ and $\Omega$ once the partition function is known.
@@
The optimal estimate of temperature represents the temperature of the finite-sized system in this measurement. In practice, by substituting the measured energy into Eqs (6) and (7), the resulting estimator corresponds to the temperature of the finite-sized system for that particular measurement according to Mandelbrot’s viewpoint [56, 58, 59, 60]. This approach provides a theoretical foundation for analyzing temperature fluctuations. In fact, that is how temperature fluctuations are measured in experiments. For example, in the paramagnetic salt HRT [21, 22, 23], temperature is obtained by measuring changes in the magnetization. In our theoretical framework, this is not optimal and can be improved by measuring energy instead.
III Estimation of temperature for large-$N$ systems
The UMVUE of $T$ and $\beta$, $\hat{T}_{G}$ and $\hat{\beta}_{B}$, depend on the expression of $\sigma(E)$. For large-$N$ systems, we approximate the expression of $\sigma(E)$ through the partition function $Z$ and its derivatives, which can be obtained through measuring thermodynamic quantities. Using the Laplace approximation (saddle point approximation) [72, 73] or the Darwin–Fowler method in Sec. 9.1 of [67] (see Appendix E), we obtain
@@ III Estimation of temperature for large-$N$ systems
In FIG. 4, the achievable ETU approach the refined ETU as $N$ increases. As $N\to\infty$, the skewness approaches 0 due to the CLT. The Gaussian-type temperature fluctuation and the usual ETU are recovered.
@@
Because $\hat{T}_{G}$ is the UMVUE of $T$, its variance is the infimum of among variances of unbiased estimators of $T$, i.e.,
@@
In the large-$N$ limit, the optimal estimation of temperature leads to a refined $\Delta T\Delta E/T^{2}$ using Eq. (15) (see Appendix E):
@@
Also, the second term on the right-hand side of Eq. (17) approaches 0 as $N\to\infty$ due to the CLT.
As an example, we consider classical ideal gas to illustrate the estimation of temperature (see Appendix F).
IV Estimation of temperature in repeated sampling
In single sampling (obtaining one outcome of the system energy), the finite-sized system exhibits significant fluctuations. To improve the precision of measurement, we discuss the UMVUE of temperature in repeated sampling with the sample size $M$. Let $\bm{i}$$\left(i_{1},i_{2},...,i_{M}\right)$ denote the microstates in repeated sampling. The $M$-independent sampling is equivalent to a single sampling of $M$-identical non-interacting systems. Here, we call the $M$-identical non-interacting systems “$M$-sampling ensemble”.

### v5

@@ Abstract
keywords:
temperature fluctuations, optimal estimation, energy-temperature uncertainty
articletype: Paper
1 Introduction
Temperature fluctuations become increasingly apparent as the system size decreases. Due to random heat transfer between a finite-sized system and its reservoir, their temperatures are not always equal. This behavior challenges the zeroth law of thermodynamics and the notion of a well-defined temperature for the finite-sized system. Experimentally, Gaussian-type temperature fluctuations have been observed in paramagnetic salt high resolution thermometer (HRT) [1, 2, 3], ferromagnetic alloy HRT [4], and nanoscale electron calorimeter [5]. It is verified that the temperature of the system exhibits fluctuations, while the temperature of the reservoir remains fixed.
Such challenges manifest in various research areas. (a) Definition of temperature in finite-sized systems including the definition of entropy and temperature in isolated systems [6, 7, 8, 9, 10], the existence of negative absolute temperatures [11, 12, 13, 14, 15, 16], the energy–temperature uncertainty relation (ETU) [17, 18] and the thermodynamics of finite-sized systems [19, 20, 21, 22, 23]; (b) Uncertainty of temperature measurement in quantum thermometry [24, 25, 29, 26, 27, 28] and nanoelectromechanical resonators [30, 31]; (c) Temperature fluctuations in nonequilibrium systems including superstatistics [32, 33, 34], nonequilibrium processes and stochastic thermodynamics [35, 36, 37, 38, 39]. Despite considerable progress, conclusions drawn from specific viewpoints or scenarios are sometimes inconsistent, and a unified understanding of thermal fluctuations in finite-sized systems remains incomplete. This situation reflects the absence of a general mathematical framework for studying such fluctuations, which motivates our work to develop a more systematic approach.
Historically, Einstein firstly investigated equilibrium fluctuations [40, 43, 41, 42] by relating the probability of macroscopic states to entropy through an inversion of the Boltzmann’s entropy formula  [^1] The operation of inverting formula of entropy has been further developed and extended to Tsallis statistics to describe nonequilibrium fluctuations of strongly correlated systems  [44].. Bohr suggested that there should exist a form of complementarity between temperature and energy in thermodynamics similar to that of position and momentum in quantum theory [17, 18]. Then, Mandelbrot linked the ETU ($\beta$ denotes the inverse temperature of the system and $E$ denotes the energy of the system)
@@ 1 Introduction
to the Cramér–Rao bound in estimation theory through the polemic exchange with Feshbach and Kittel [45, 46, 48, 47], where $\Delta(\cdot)=\sqrt{\langle\cdot^{2}\rangle-\langle\cdot\rangle^{2}}$. This practice originated from 1956 [49] when Mandelbrot applied the estimation theory to the derivation of statistical thermodynamics from purely phenomenological principles (also see [50, 51, 52]). It inspires us to use the uncertainty of temperature measurement to understand temperature fluctuations in finite-sized systems. For more dicussions about thermodynamic uncertainty relations, we refer to [53].
@@
In repeated sampling, the optimal estimation of temperature exhibits sample-size dependence, coinciding with their counterparts in nanothermodynamics [19, 20, 21, 22, 23]. Actually, the “sampling ensemble” can be regarded as the statistical interpretation of the replica trick in Hill’s nanothermodynamics. And a Gaussian-type temperature fluctuation naturally emerges from a general temperature distribution due to the central limit theorem (CLT).
Compared with quantum thermometry, which mainly focuses on improving the precision of temperature measurements (especially for low temperatures) through dynamical [24, 25, 26, 27] or steady-state properties[29, 30, 27, 31, 28], we aim to describe temperature fluctuations in finite-sized systems through a rigorous mathematical framework. Therefore, unlike thermometry, which typically seeks to improve precision by modifying system’s energy spectrum, optimizing probes, or performing sufficiently many measurements, we focus on the optimal estimation of temperature for a given energy spectrum with a finite number of measurements. Moreover, the UMVUE derived in this work provides an achievable bound that is tighter than the Cramér–Rao bound, potentially shedding light on the generation of more accurate methods of temperature estimation.
2 Estimation of temperature in single sampling
Developing Mandelbrot’s viewpoint about temperature fluctuations [48, 49, 50, 51], i.e., “temperature for systems-in-isolation should be viewed as a statistical estimate of the parameter of a conjectural canonical distribution”, we treat a finite-sized system in canonical ensemble as a thermometer of a reservoir. The $N$-particle system contacts a real or hypothetical heat reservoir with temperature $T$. At equilibrium, the probability of the system at state $i$ obeys the Boltzmann-Gibbs formula:
@@ 2 Estimation of temperature in single sampling
where $\beta=T^{-1}$ denotes the inverse temperature, $E_{i}$ $(d_{i})$ denotes the energy (degeneracy) of energy level with label $i\in D$, $D$ denotes the set of energy level indices, and $Z=\sum_{i\in D}d_{i}e^{-\beta E_{i}}$ is the partition function.
@@
where $\sigma(E)\equiv\sum_{i\in D}\delta(E-E_{i})d_{i}$ denotes the density of states and  [^2] Eq. (7) was noted by Gibbs [56], but its connection to estimation theory was not explored.
$\text{d}^{k}_{E}\sigma(E)\equiv\begin{cases}\frac{\text{d}^{k}\sigma(E)}{\text{d}E^{k}}&\text{for }k\textgreater 0\\ \sigma(E)&\text{for }k=0\\ \int_{E_{\text{min}}}^{E}\text{d}E^{{}^{\prime}}\text{d}_{E^{\prime}}^{k+1}\sigma(E^{\prime})&\text{for }k\textless 0\end{cases}.$ (7)
For $k=1$, we write $\text{d}^{1}_{E}\sigma(E)$ as $\text{d}_{E}\sigma(E)$ for simplicity. In the derivation, we have assumed the density of states $\sigma(E)$ satisfies
@@
for integer $l\in[0,k)$, which holds in classical systems but not in quantum systems. This is relevant to the third law of thermodynamics  [^3] The third law of thermodynamics is a macroscopic manifestation of quantum effects [57, 58, 59]. One formulation of the third law is $\lim_{T\to 0}S=0$, which corresponds to the system’s energy being $E_{\min}$. If $\sigma(E_{\min})=1$, the law holds; if $\sigma(E_{\min})=0$, the limit $\lim_{T\to 0}S$ does not exist, and the law fails..
When the above assumption (8) is invalid, the estimator of $\beta^{k}$ for $k<0$, as well as that of $\beta$ given in Eq. (6) are no longer unbiased and instead exhibit an exponentially decaying bias with $N$. In figure 2, we illustrate this bias in an analogous low-temperature Fermi gas system using relative bias $\Phi_{\beta}=|\langle\hat{\beta}^{(1)}\rangle-\beta|/|\beta|$. Therefore, when $N$ is sufficiently large, the bias becomes negligible, and Eq. (6) still provides their unbiased estimators.
@@
It is worth noting that $k$ imposed by the above estimators have an upper bound (in this paper we mainly consider the case $k=\pm 1$), depending on the degrees of freedom of the system. As a example, taking the density of states $\sigma(E)=E^{3N/2-1}$ of a classical ideal gas as an example [57], $\hat{\beta}^{(k)}$ in Eq. (6) is
@@
where $k\leq 3N/2-2$ and $\Gamma$ denotes the Gamma function  [^4] When $k\geq 3N/2-1$, the proof in Appendix A used to verify that $\hat{\beta}^{(k)}$ is an unbiased estimator of $\beta^{k}$ no longer holds..
@@
Secondly, when $E_{\text{max}}$ is finite, the estimators of the powers of inverse temperature $(k>0)$ given in Eq. (6) are no longer unbiased and instead exhibit a bias that decays exponentially with $N$. In figure 2, we illustrate this bias in an $N$ non-interacting two-level systems, showing that it leads to a relative bias $\Phi_{\beta}\sim\text{e}^{-\alpha_{1}N}$, where the decay rate $\alpha_{1}=\ln(1+\text{e}^{-|\beta|\varepsilon})$ (see Appendix C), and $\varepsilon$ denotes the spacing of energy levels. Therefore, when $N$ is sufficiently large, the bias becomes negligible, and Eq. (6) still provides their unbiased estimators.
For the estimators of the powers of temperature $(k<0)$, we have to consider the cases separately. For positive temperatures, the estimators given by Eq. (6) exhibit a analogous bias that decays exponentially with $N$, which becomes negligible once $N$ is sufficiently large. In contrast, for negative temperatures, using Eq. (6) to estimate the system’s temperature results in a bias that grows exponentially with $N$, which demonstrates that Gibbs temperature cannot be used to estimate the temperature of a negative-temperature system. In figure 3, we illustrate these biases in an $N$ non-interacting two-level systems, showing that they lead to a relative bias $\Phi_{T}=|\langle\hat{\beta}^{(-1)}\rangle-T|/|T|\sim\text{e}^{-\alpha_{2}N}$, where the decay (increase) rate $\alpha_{2}=\frac{\beta\varepsilon}{2}+\ln\cosh(\frac{\beta\varepsilon}{2})$ (see Appendix C). The above indicates that an optimal estimator for the temperature of a negative-temperature system likely does not exist. Moreover, we extend the unbiased estimator of $\beta^{\alpha}$ to $\alpha\in\mathbb{R}$ (see Appendix D).
@@
A corollary following the RBLS theorem [54, 61, 62] is: If $T$ is a complete sufficient statistic for the parameter $\theta$, then an estimator $\tilde{\theta}(T)$ that depends only on $T$ is the UMVUE of its expected value.
The Boltzmann–Gibbs formula in Eq. (5) belongs to the exponential family in statistical inference, and its complete sufficient statistic is $E_{i}$ [54]. According to the above corollary, since the estimators $\hat{\beta}^{(k)}$ in Eq. (6) are not only unbiased but also a functions of $E_{i}$, they are the UMVUEs of $\beta^{k}$. Moreover, no simple functional relationship exists among the UMVUEs of different moments of $\beta$ implies that knowing the optimal estimator of the temperature does not directly yield the estimator of any function of the temperature.
Figure 3: Relative bias of $T$ varies as NN when $E_{\text{max}}$ is finite. Here, the density of states is the same as figure 2 (see Appendix C). (a) $T>0$. (b) $T<0$.
@@
Hence, the Boltzmann entropy $S_{B}$ and the Gibbs entropy $S_{G}$ correspond to the UMVUE of inverse temperature and temperature respectively. It is worth mentioning that $\hat{\beta}_{B}$ and $\hat{T}_{G}$ have been discussed in previous studies. However, there are various controversies among these studies, mainly concerning the existence of negative temperatures [11, 12, 13, 14, 15, 16], the validity of the equipartition theorem [11, 9, 10], the requirements of the zeroth law of thermodynamics [7, 13], the additivity of entropy, adiabatic invariance, and more [7, 8, 9]. These controversies arise because most studies start from the four laws of thermodynamics and empirical facts valid in the thermodynamic limit. In contrast, we base our approach on the mathematical theorems in statistical inference, in which different estimators correspond to different formulas of entropy and temperature.
@@
    Determine the expressions for $\sigma$ and $\Omega$, which are either known before the measurement or can be obtained using the method in Section 3. This method relates $\sigma$ to the partition function and its derivatives in the large-$N$ limit, allowing us to calculate $\sigma$ and $\Omega$ once the partition function is known.
@@
The optimal estimate of temperature represents the temperature of the finite-sized system in this measurement, following Mandelbrot’s viewpoint [48, 49, 50, 51]. In quantum thermometry, it has been demonstrated that energy measurements are the most informative measurement [27]. It is worth noting that temperature fluctuations may also arise from reservoirs with finite capacity (see Ref. [39] for details).
3 Estimation of temperature for large-$N$ systems
The UMVUE of $T$ and $\beta$, $\hat{T}_{G}$ and $\hat{\beta}_{B}$, depend on the expression of $\sigma(E)$. For large-$N$ systems, we approximate the expression of $\sigma(E)$ through the partition function $Z$ and its derivatives, which can be obtained through measuring heat capacity of the system. Using the Laplace approximation (saddle point approximation) [63, 64] or the Darwin–Fowler method in Sec. 9.1 of [57] (see Appendix E), we obtain
@@ 3 Estimation of temperature for large-$N$ systems
In figure 4, the achievable ETU approach the refined ETU as $N$ increases. As $N\to\infty$, the skewness approaches 0 due to the CLT. The Gaussian-type temperature fluctuation and the usual ETU are recovered.
@@
And we have
@@
The optimal estimation of temperature leads to a refined $\Delta T\Delta E/T^{2}$ in the large-$N$ limit (see Appendix E):
@@
As an illustration, the estimation of temperature in classical ideal gas is considered (see Appendix F).
4 Estimation of temperature in repeated sampling
In single sampling, the measured energy of the finite-sized system exhibits significant fluctuations. To improve the precision of measurement, we discuss the UMVUE of temperature in repeated sampling with the sample size $M$. Let $\bm{i}$$\left(i_{1},i_{2},...,i_{M}\right)$ denote the microstates in repeated sampling. The $M$-independent sampling is equivalent to a single sampling of $M$-identical non-interacting systems. Here, we call the $M$-identical non-interacting systems “$M$-sampling ensemble”.